# An Ising Similarity Regression Model

### v1

@@ 2.2 Estimation and Selection
Suppose we have observations of $p$-dimensional response vectors $\bm{y}_{i}=(y_{i1},\cdots,\\ y_{ip})^{\top}\in\{0,1\}^{p}$ for $i=1,\cdots,n$, where $\bm{y}_{1},\cdots,\bm{y}_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}f(\cdot;\bm{\vartheta})$ follows the Ising similarity regression model in equation (2.4), and $i.i.d$ denotes independent and identically distributed. Recall the pmf $f(\cdot;\bm{\vartheta})$ involves the normalization constant $Z(\bm{\vartheta})=\sum_{\bm{u}\in\{0,1\}^{p}}\exp\{\sum_{j=1}^{p}\theta_{jj}u_{j}+\sum_{1\leq j<j^{\prime}\leq p}(\sum_{k=1}^{K}\alpha_{k}w_{jj^{\prime}}^{(k)})u_{j}u_{j^{\prime}}\}$, which is a sum of $2^{p}$ terms. As a result, maximum likelihood estimation based on $\prod_{i=1}^{n}f(\bm{y}_{i};\bm{\vartheta})$ is computationally not feasible when the dimension $p$ is large. To overcome this problem, we adapt the existing literature (e.g., Höfling and Tibshirani, 2009; Ravikumar, Wainwright, and Lafferty, 2010) and propose a pseudo-likelihood estimation approach to fit the Ising similarity regression model. Let $\bm{y}_{i\backslash j}=(y_{i1},\cdots,y_{i,j-1},y_{i,j+1},\cdots,y_{ip})^{\top}$ denote the $i$-th observed response vector without the $j$-th element, and $f_{j}(\cdot|\bm{y}_{i\backslash j};\bm{\vartheta})$ denote the corresponding conditional pmf of $y_{ij}$ given $\bm{y}_{i\backslash j}$ for $i=1,\cdots,n$ and $j=1,\cdots,p$. Then the (unregularized) pseudo-likelihood estimator can be obtained by maximizing the pseudo-likelihood function $\prod_{i=1}^{n}\prod_{j=1}^{p}f_{j}(y_{ij}|\bm{y}_{i\backslash j};\bm{\vartheta})$. In particular, the conditional pmf $f_{j}(\cdot|\bm{y}_{i\backslash j};\bm{\vartheta})$ can be derived based on $f(\cdot;\bm{\vartheta})$ in equation (2.4), and takes the simple form below,
$f_{j}(u|\bm{y}_{i\backslash j};\bm{\vartheta})$ =$=\{\mathrm{P}(y_{ij}=1|\bm{y}_{i\backslash j};\bm{\vartheta})\}^{u}\{1-\mathrm{P}(y_{ij}=1|\bm{y}_{i\backslash j};\bm{\vartheta})\}^{1-u}$
@@
$=\frac{\exp\left\{u\left(\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\sum_{j^{\prime}\neq j}w_{jj^{\prime}}^{(k)}y_{ij^{\prime}}\right)\right\}}{1+\exp\left(\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\sum_{j^{\prime}\neq j}w_{jj^{\prime}}^{(k)}y_{ij^{\prime}}\right)},$ (2.5)
@@
We can further augment the above pseudo-likelihood estimator with a penalty to perform variable selection on the elements of $\bm{\alpha}$. This is useful in practice when there are a non-negligible number of similarity matrices available for data analysis, but only a subset of them are anticipated to be relevant in equation (2.3). For example, our real application in Section 5 considered 15 similarity matrices for the U.S. senators’ voting records data, but only a subset of these were selected to be informative for driving conditional dependence relationships in their voting patterns.
To perform simultaneous estimation and regularization on the coefficient vector $\bm{\alpha}$, we augment the log pseudo-likelihood function based on equation (2.5) with an adaptive lasso penalty (Zou, 2006), resulting in a regularized pseudo-likelihood estimator that minimizes the objective function
@@
given a tuning parameter $\lambda>0$. Following Zou (2006) and Huang, Ma, and Zhang (2008) among others, we set the adaptive weights as $w_{k}=1/|\bar{\alpha}_{k}|$ for $k=1,\cdots,K$, where $\bar{\bm{\alpha}}=(\bar{\alpha}_{1},\cdots,\bar{\alpha}_{K})^{\top}$ denotes the unregularized pseudo-likelihood estimator i.e., the estimator of $\bm{\alpha}$ which minimizes the objective function (2.7) with $\lambda=0$. It is worth noting that $\theta_{jj}$, the $j$-specific intercepts for $j=1,\cdots,p$, are not regularized; this is similar to the literature of sparse Ising models (Guo et al., 2010; Ravikumar, Wainwright, and Lafferty, 2010) as well as other regularized regression settings in general (e.g, Hui, Müller, and Welsh, 2017b, 2018). Regarding the choice of the penalty function, in this article we focus our developments on the adaptive lasso penalty $\lambda\sum_{k=1}^{K}w_{k}\left|\alpha_{k}\right|$; see Section 6 for a discussion on alternative penalty functions. In particular, with adaptive weights there is only a single regularization parameter $\lambda$ in the penalty, and the whole objective function (2.7) remains convex for optimization. Moreover, the incorporation of adaptive weights $w_{k}$ allows for varying degrees of regularization on the coefficients $\alpha_{k}$ across $k=1,\cdots,K$, and facilitates selection consistency for a sparse coefficient vector $\bm{\alpha}$ which we theoretically examine in the next section. The details of implementing the regularized pseudo-likelihood estimation by minimizing (2.7) are discussed in Section 4.
@@ 3 Theoretical Results
$\hat{\bm{\alpha}}=\min_{\bm{\alpha}}\left\{-l(\bm{\alpha})+\lambda\sum_{k=1}^{K}w_{k}\left|\alpha_{k}\right|\right\},$ (3.8)
where $l(\bm{\alpha})=\sum_{i=1}^{n}l_{i}(\bm{\alpha})/(np)$ and $l_{i}(\bm{\alpha})=\sum_{j=1}^{p}[y_{ij}(\sum_{k=1}^{K}\alpha_{k}\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i})-\log\{1+\exp(\sum_{k=1}^{K}\alpha_{k}$ $\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i})\}]$; see Guo et al. (2010, 2015) for a similar theoretical treatment. As discussed previously, the adaptive weights $w_{k}=1/|\bar{\alpha}_{k}|$ in criterion (3.8) are constructed based on the elements in the vector of the unregularized pseudo-likelihood estimator $\bar{\bm{\alpha}}=\arg\min_{\bm{\alpha}}\{-l(\bm{\alpha})\}$. With this in mind, we will first investigate the asymptotic properties of $\bar{\bm{\alpha}}$.
@@
and define the $K\times K$ matrices $\bm{U}^{0}=E\{\sum_{i=1}^{n}\bm{\mathcal{X}}^{(i)\top}\bm{\mathcal{X}}^{(i)}/(np)\}$ and $\bm{M}^{0}=E\{-\nabla^{2}l(\bm{\alpha}^{(0)})\}$. Finally, for a generic $m\times r$ matrix $\bm{H}=(h_{tv})_{m\times r}$, and subsets of row and column indices $\mathcal{T}\subseteq\{1,\cdots,m\}$ and $\mathcal{V}\subseteq\{1,\cdots,r\}$, let $\bm{H}_{\mathcal{T},\mathcal{V}}$ denote the submatrix of $\bm{H}$ consisting of rows and columns indexed by $\mathcal{T}$ and $\mathcal{V}$, respectively, $\left\|\bm{H}\right\|_{\infty}=\max_{1\leq t\leq m}\{\sum_{v=1}^{r}|h_{tv}|\}$ denote the matrix infinity-norm of $\bm{H}$, and $\Lambda_{\min}(\bm{H})$ and $\Lambda_{\max}(\bm{H})$ denote the smallest and largest eigenvalues of $\bm{H}$, respectively, when $\bm{H}$ is a square matrix (i.e., $m=r$). We introduce the following technical conditions.
@@
There exist finite positive constants $C_{\min}$ and $C_{\max}$ such that $\Lambda_{\min}(\bm{M}^{0})\\ \geq C_{\min}$ and $\Lambda_{\max}(\bm{U}^{0})\leq C_{\max}$.
@@
There exists a finite positive constant $C_{W}$ such that $\left\|\bm{W}_{k}\right\|_{\infty}\leq C_{W}$ for all symmetric similarity matrices $\{\bm{W}_{k}:k=1,\cdots,K\}$.
By the Cauchy Interlacing Theorem (Parlett, 1980, p. 186) and Condition 1, similar conditions hold for the submatrices $\bm{M}_{S,S}^{0}$ and $\bm{U}_{S,S}^{0}$; in particular, $\Lambda_{\min}(\bm{M}_{S,S}^{0})\geq C_{\min}$ and $\Lambda_{\max}(\bm{U}_{S,S}^{0})\leq C_{\max}$. Condition 2 is similar to the conditions imposed on the similarity matrices $\bm{W}_{k}$ in the literature (see for instance Zou et al., 2017). We first establish the estimation consistency for the unregularized pseudo-likelihood estimator.
@@
Assume Conditions 12 are satisfied. If $K\sqrt{\log(p)/n}=o(1)$ and there exists a finite positive constant $C_{\nabla}$ such that $K=o(p^{C_{\nabla}^{2}/(8C_{W}^{2})})$ as $n,p\to\infty$, then with probability tending to one it holds that
$\left\|\bar{\bm{\alpha}}-\bm{\alpha}^{(0)}\right\|_{2}\leq\bar{M}\sqrt{\frac{K\log(p)}{n}},$
for a finite positive constant $\bar{M}>4C_{\nabla}/C_{\min}$.
@@
To conclude this section, we remark Condition 4 provides the requirements on the rate of the tuning parameter $\lambda$ based on the smallest truly non-zero coefficient $\min\{|\alpha^{(0)}_{k}|:k\in S\}$, as the basis for establishing the results in Theorem 2. For instance, if $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ is bounded away from zero, then it can be verified that $\lambda=\{n/\log(p)\}^{-t}K^{v}$ for some $t\in(1/2,1)$ and $v\in[2t-3/2,2t-1]$ suffices for Condition 4 to hold. If this is relaxed and we permit $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ to tend to zero at a rate satisfying $\min\{|\alpha^{(0)}_{k}|:k\in S\}\{n/\log(p)\}^{m}\to\infty$ for some $m\in(0,1/4]$, then $\lambda=O(\{\log(p)/n\}^{q})$ for some $q\in[m+1/2,3/4]$ will satisfy the requirements. In practice, since $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ is unknown, then we adopt a data-driven approach to select $\lambda$ as discussed in the next section.
@@ 4 Simulation Study
We conducted a numerical study to evaluate the finite sample performance of the regularized pseudo-likelihood estimator for the Ising similarity regression model (2.4). Following Guo et al. (2010), we generated the true main effect parameters $\{\theta_{jj}^{(0)};j=1,\cdots,p\}$ independently and uniformly in the range $[-1,-0.5]\cup[0.5,1]$. For the regression coefficients of the similarity matrices, we set $K=20$ and generated $\alpha^{(0)}_{k}$ independently and uniformly from $[-0.4,-0.3]\cup[0.3,0.4]$ for $k=1,\cdots,5$ and set $\alpha^{(0)}_{k}=0$ for $k=6,\cdots,20$. This led to $K_{0}=5$ truly non-zero coefficients. Turning to the similarity matrices, we followed the framework of Zou et al. (2017) and set the diagonal elements of all the similarity matrices $\{\bm{W}_{k};k=1,\cdots,K\}$ to zero, while the off-diagonal elements were simulated as $w_{jj^{\prime}}^{(k)}=w_{j^{\prime}j}^{(k)}=\exp(-d_{jj^{\prime}}^{(k)2})$ with $d_{jj^{\prime}}^{(k)}\stackrel{{\scriptstyle i.i.d}}{{\sim}}U(p^{-1/2},p^{1/2})$ for $j,j^{\prime}=1,\cdots,p$. Based on the above settings, we then constructed the true Ising similarity regression model as in (2.4), and afterward simulated datasets consisting of $n$ multivariate binary vectors of dimension $p$ using the R package IsingSampler (Epskamp, 2020). We considered sample sizes $n\in\{50,100,200,400\}$ and different numbers of binary responses $p\in\{10,25,50,100,200\}$, and simulated 1000 datasets for each combination of $n$ and $p$.
@@
We employed a two-step algorithm to compute the regularized pseudo-likelihood estimator. First, from the discussion below equations (2.6) – (2.7), the case of $\lambda=0$ i.e., the unregularized pseudo-likelihood estimator, can be obtained by fitting a logistic regression model with $\bm{y}_{1:n}=(y_{11},\cdots,y_{n1},\cdots,y_{1p},\cdots,y_{np})^{\top}$ as the response and $(\bm{I}_{p}\otimes\bm{1}_{n},\bm{\mathcal{X}})$ as the model matrix, where $\bm{1}_{n}$ is an $n$-dimensional vector of ones, $\otimes$ is the Kronecker product operator, and
$\bm{\mathcal{X}}=\begin{pmatrix}\bm{W}_{1\cdot}^{(1)\top}\bm{y}_{1}&\cdots&\bm{W}_{1\cdot}^{(K)\top}\bm{y}_{1}\\ \vdots&\vdots&\vdots\\ \bm{W}_{1\cdot}^{(1)\top}\bm{y}_{n}&\cdots&\bm{W}_{1\cdot}^{(K)\top}\bm{y}_{n}\\ \vdots&\vdots&\vdots\\ \bm{W}_{p\cdot}^{(1)\top}\bm{y}_{1}&\cdots&\bm{W}_{p\cdot}^{(K)\top}\bm{y}_{1}\\ \vdots&\vdots&\vdots\\ \bm{W}_{p\cdot}^{(1)\top}\bm{y}_{n}&\cdots&\bm{W}_{p\cdot}^{(K)\top}\bm{y}_{n}\\ \end{pmatrix}.$
After obtaining the unregularized estimator $\bar{\bm{\alpha}}$ in the first step, we used its elements $\bar{\alpha}_{k}$’s to construct adaptive regularization weights in (2.7) before proceeding to compute the regularized pseudo-likelihood estimator by fitting an adaptive lasso regularized logistic regression to $\bm{y}_{1:n}$ and $(\bm{I}_{p}\otimes\bm{1}_{n},\bm{\mathcal{X}})$ in the second step. We used the R package glmnet (Friedman, Hastie, and Tibshirani, 2010) to accomplish this. To select the tuning parameter $\lambda$ in (2.7), we utilized a ten-fold cross-validation approach, where the observations in each dataset were grouped at the level indexed by $i=1,\cdots,n$. That is, we split the data $\{(\bm{y}_{1},\bm{\mathcal{X}}^{(1)}),\cdots,(\bm{y}_{n},\bm{\mathcal{X}}^{(n)})\}$ randomly into ten folds, fitted the regularized logistic regression to nine of the (training) folds, and then computed the pseudo-likelihood of the fitted model on the remaining (test) fold. This was then repeated so that each fold acted as a test set once, noting that the same set of adaptive regularization weights constructed based on the unregularized estimator from the full dataset was used throughout. We then selected an optimal $\lambda$ that gave the largest mean pseudo-likelihood (averaged across the ten test sets). Note the grouping of the observations was done to preserve the dependence structure of the data, since $y_{ij}$ were independent in the index $i$ but dependent in the response index $j$ through the Ising model (see also Warton, Thibaut, and Wang, 2017). We compared the performance of our proposed regularized pseudo-likelihood estimator against two other basic approaches: the unregularized estimator based on (2.7) with $\lambda=0$ (i.e., no regularization), and the oracle estimator based on minimizing an alternative to (2.7) where $\lambda=0$ and $\alpha_{k}$ are set to be zeros for $k=6,\cdots,20$. The unregularized estimator was included to examine whether regularization could lead to better overall estimation performance for the regression coefficients and main effect parameters, while the oracle estimator was used as a benchmark since it incorporated additional information regarding the index set $S=\{1,\cdots,5\}$. For all three methods compared, we evaluated the performance of the point estimates for the regression coefficients $\alpha_{k}$ and main effect parameters $\theta_{jj}$ using the mean square error (MSE), $\textrm{MSE}_{\alpha}=(1/1000)\sum_{l=1}^{1000}\sum_{k=1}^{20}(\hat{\alpha}_{k,\{l\}}-\alpha^{(0)}_{k})^{2}/K$ and $\textrm{MSE}_{\theta}=(1/1000)\sum_{l=1}^{1000}\sum_{j=1}^{p}\\ (\hat{\theta}_{jj,\{l\}}-\theta^{(0)}_{jj})^{2}/p$, respectively, noting the MSE was scaled by the number of associated parameters, and $\hat{\alpha}_{k,\{l\}}$ and $\hat{\theta}_{jj,\{l\}}$ generically denote the estimated regression coefficients and main effect parameters, respectively, from the $l$-th simulated dataset. Additionally, we assessed model selection performance for the regularized estimators of the regression coefficients based on the true positive rate (TPR) and false positive rate (FPR), where $\textrm{TPR}=(1/1000)\sum_{l=1}^{1000}\sum_{k=1}^{5}\\ \bm{1}(\hat{\alpha}_{k,\{l\}}\neq 0)/5$, $\textrm{FPR}=(1/1000)\sum_{l=1}^{1000}\sum_{k=6}^{20}\bm{1}(\hat{\alpha}_{k,\{l\}}\neq 0)/15$, and $\bm{1}(\cdot)$ is the indicator function.

### v2

@@ 2.2 Estimation and Similarity Selection
Suppose we have observations of $p$-dimensional response vectors $\bm{y}_{i}=(y_{i1},\cdots,\\ y_{ip})^{\top}\in\{0,1\}^{p}$ for $i=1,\cdots,n$, where $\bm{y}_{1},\cdots,\bm{y}_{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}f(\cdot;\bm{\vartheta})$ follows the Ising similarity regression model in equation (2.4), and $i.i.d.$ denotes independent and identically distributed. Recall the pmf $f(\cdot;\bm{\vartheta})$ involves the normalization constant $Z(\bm{\vartheta})$, which is a sum of $2^{p}$ terms. As a result, maximum likelihood estimation based on $\prod_{i=1}^{n}f(\bm{y}_{i};\bm{\vartheta})$ is computationally not feasible when the dimension $p$ is large. To overcome this problem, we adapt the existing literature (e.g., Höfling and Tibshirani, 2009; Ravikumar, Wainwright, and Lafferty, 2010) and propose a pseudo-likelihood estimation approach to fit the Ising similarity regression model. Let $\bm{y}_{i\backslash j}=(y_{i1},\cdots,y_{i,j-1},y_{i,j+1},\cdots,y_{ip})^{\top}$ denote the $i$-th observed response vector without the $j$-th element, and $f_{j}(\cdot|\bm{y}_{i\backslash j};\bm{\vartheta})$ denote the corresponding conditional pmf of $y_{ij}$ given $\bm{y}_{i\backslash j}$ for $i=1,\cdots,n$ and $j=1,\cdots,p$. Then the (unregularized) pseudo-likelihood estimator can be obtained by maximizing the pseudo-likelihood function $\prod_{i=1}^{n}\prod_{j=1}^{p}f_{j}(y_{ij}|\bm{y}_{i\backslash j};\bm{\vartheta})$. In particular, the conditional pmf $f_{j}(\cdot|\bm{y}_{i\backslash j};\bm{\vartheta})$ can be derived based on $f(\cdot;\bm{\vartheta})$ in equation (2.4), and takes the simple form below,
@@
$f_{j}(u|\bm{y}_{i\backslash j};\bm{\vartheta})=\frac{\exp\left\{u\left(\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\sum_{j^{\prime}\neq j}w_{jj^{\prime}}^{(k)}y_{ij^{\prime}}\right)\right\}}{1+\exp\left(\theta_{jj}+\sum_{k=1}^{K}\alpha_{k}\sum_{j^{\prime}\neq j}w_{jj^{\prime}}^{(k)}y_{ij^{\prime}}\right)},$ (2.5)
@@
We can further augment the above pseudo-likelihood estimator with a penalty to perform variable selection on the elements of $\bm{\alpha}$. This is useful in practice when there are a non-negligible number of similarity matrices available for data analysis, but only a subset of them are anticipated to be relevant in equation (2.3). To perform simultaneous estimation and regularization on the coefficient vector $\bm{\alpha}$, we augment the log pseudo-likelihood function based on equation (2.5) with an adaptive lasso penalty (Zou, 2006), resulting in a regularized pseudo-likelihood estimator that minimizes the objective function
@@
given a tuning parameter $\lambda>0$. Following Zou (2006) and Huang, Ma, and Zhang (2008) among others, we set the adaptive weights as $w_{k}=1/|\bar{\alpha}_{k}|$ for $k=1,\cdots,K$, where $\bar{\bm{\alpha}}=(\bar{\alpha}_{1},\cdots,\bar{\alpha}_{K})^{\top}$ denotes the unregularized pseudo-likelihood estimator i.e., the estimator of $\bm{\alpha}$ which minimizes the objective function (2.7) with $\lambda=0$. It is worth noting that $\theta_{jj}$, the $j$-specific intercepts for $j=1,\cdots,p$, are not regularized; this is similar to the literature of sparse Ising models (Guo et al., 2010; Ravikumar, Wainwright, and Lafferty, 2010) as well as other regularized regression settings in general (e.g, Hui, Müller, and Welsh, 2017b, 2018). Regarding the choice of the penalty function, in this article we focus our developments on the adaptive lasso penalty $\lambda\sum_{k=1}^{K}w_{k}\left|\alpha_{k}\right|$; see Section 6 for a discussion on alternative penalty functions. In particular, with adaptive weights there is only a single regularization parameter $\lambda$ in the penalty, and the whole objective function (2.7) remains convex for optimization. Moreover, the incorporation of adaptive weights $w_{k}$ allows for varying degrees of regularization on the coefficients $\alpha_{k}$ across $k=1,\cdots,K$, and facilitates selection consistency for a sparse coefficient vector $\bm{\alpha}$ which we theoretically examine in the next section. We also emphasize that the adaptive lasso penalty here is used for the selection of regression coefficients associated with similarity matrices i.e., similarity selection, and not for the selection of edges in $\bm{\Theta}$. This is apparent when we see that the penalty induces sparsity in the regression coefficients, but not the similarity matrices and the resulting $\bm{\Theta}$. The details of implementing the regularized pseudo-likelihood estimation by minimizing (2.7) are discussed in Section 4.
@@ 3 Theoretical Results
$\hat{\bm{\alpha}}=\arg\min_{\bm{\alpha}}\left\{-l(\bm{\alpha})+\lambda\sum_{k=1}^{K}w_{k}\left|\alpha_{k}\right|\right\},$ (3.8)
where $l(\bm{\alpha})=\sum_{i=1}^{n}l_{i}(\bm{\alpha})/(np)$, $l_{i}(\bm{\alpha})=\sum_{j=1}^{p}[y_{ij}(\sum_{k=1}^{K}\alpha_{k}\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i})-\log\{1+\exp(\sum_{k=1}^{K}\alpha_{k}$ $\bm{W}_{j\cdot}^{(k)\top}\bm{y}_{i})\}]$, $w_{k}=1/|\bar{\alpha}_{k}|$ and $\bar{\bm{\alpha}}=(\bar{\alpha}_{1},\cdots,\bar{\alpha}_{K})^{\top}=\arg\min_{\bm{\alpha}}\{\\ -l(\bm{\alpha})\}$; see Guo et al. (2010, 2015) for a similar theoretical treatment. It is worth noting that while the pseudo-likelihood function is used in criterion (3.8) to overcome the issue of intractable normalization constant, theoretical results in this section are still obtained under the joint probability distribution of the Ising similarity regression model.
@@
and define the $K\times K$ matrices $\bm{U}^{0}=E\{\sum_{i=1}^{n}\bm{\mathcal{X}}^{(i)\top}\bm{\mathcal{X}}^{(i)}/(np)\}$ and $\bm{M}^{0}=E\{-\nabla^{2}l(\bm{\alpha}^{(0)})\}$. Finally, for a generic $m\times r$ matrix $\bm{H}=(h_{tv})_{m\times r}$, and subsets of row and column indices $\mathcal{T}\subseteq\{1,\cdots,m\}$ and $\mathcal{V}\subseteq\{1,\cdots,r\}$, let $\bm{H}_{\mathcal{T},\mathcal{V}}$ denote the submatrix of $\bm{H}$ consisting of rows and columns indexed by $\mathcal{T}$ and $\mathcal{V}$, respectively, $\left\|\bm{H}\right\|_{1}=\max_{1\leq v\leq r}\{\sum_{t=1}^{m}|h_{tv}|\}$ denote the matrix 1-norm of $\bm{H}$, and $\Lambda_{\min}(\bm{H})$ and $\Lambda_{\max}(\bm{H})$ denote the smallest and largest eigenvalues of $\bm{H}$, respectively, when $\bm{H}$ is a square matrix (i.e., $m=r$). We also refer the reader to Table LABEL:table:notation in Section LABEL:sec:notation of the supplementary material for a list of important notations used throughout this article along with their definitions.
We introduce the following technical conditions.
@@
There exist finite positive constants $C_{\min}$ and $C_{\max}$ such that $\Lambda_{\min}(\bm{M}^{0})\\ \geq C_{\min}$ and $\Lambda_{\max}(\bm{U}^{0})\leq C_{\max}$, where $\bm{M}^{0}$ and $\bm{U}^{0}$ are defined below equation (3.9).
@@
There exists a finite positive constant $C_{W}$ such that $\left\|\bm{W}_{k}\right\|_{1}\leq C_{W}$ for all symmetric similarity matrices $\{\bm{W}_{k}:k=1,\cdots,K\}$.
By the Cauchy Interlacing Theorem (Parlett, 1980, p. 186) and Condition 1, similar conditions hold for the submatrices $\bm{M}_{S,S}^{0}$ and $\bm{U}_{S,S}^{0}$; in particular, $\Lambda_{\min}(\bm{M}_{S,S}^{0})\geq C_{\min}$ and $\Lambda_{\max}(\bm{U}_{S,S}^{0})\leq C_{\max}$. Condition 2 is a bounded matrix norm assumption which is similar to the conditions imposed on the similarity matrices $\bm{W}_{k}$ in the literature (see for instance, Condition C8 in Zou et al., 2022). For example, when $\bm{W}_{k}$ is a symmetric adjacency matrix with elements $w_{jj^{\prime}}^{(k)}\in\{0,1\}$ capturing the neighborhood relationship among $p$ nodes, the condition is equivalent to a column-sparsity (and row-sparsity since $\bm{W}_{k}$ are symmetric) condition that requires the number of neighbors for each node to be finite even when the total number of nodes $p$ diverges. Other classes of similarity matrices such as those with bounded elements $|w_{jj^{\prime}}^{(k)}|\leq C$ and $s$-sparse column (and row) vectors (e.g., Wainwright, 2019, p. 156) i.e., $\sum_{j=1}^{p}1_{\{|w_{jj^{\prime}}^{(k)}|>0\}}\leq s$, where $C$ and $s$ are finite positive constants and $1_{\{\cdot\}}$ is the indicator function, also satisfy Condition 2 with $C_{W}=sC$. Together with the above assumed sparsity on the true coefficient vector $\bm{\alpha}^{(0)}$, Condition 2 implies similar bounded matrix norm condition for the true interaction matrix $\bm{\Theta}^{(0)}=\sum_{k=1}^{K}\alpha^{(0)}_{k}\bm{W}_{k}$; that is, $\|\bm{\Theta}^{(0)}\|_{1}\leq C_{W}\sum_{k\in S}|\alpha^{(0)}_{k}|$.
We first establish the estimation consistency for the unregularized pseudo-likelihood estimator that is used to construct the adaptive weights.
@@
Assume Conditions 12 are satisfied. If $K\sqrt{\log(p)/n}=o(1)$ and there exists a finite positive constant $C_{\nabla}$ such that $K=o(p^{C_{\nabla}^{2}/(8C_{W}^{2})})$ as $n,p\to\infty$, then with probability tending to one it holds that $\|\bar{\bm{\alpha}}-\bm{\alpha}^{(0)}\|_{2}\leq\bar{M}\sqrt{K\log(p)/n},$ for a finite positive constant $\bar{M}>4C_{\nabla}/C_{\min}$.
@@
To conclude this section, we remark Condition 4 provides requirements on the rate of the tuning parameter $\lambda$ based on the smallest truly non-zero coefficient $\min\{|\alpha^{(0)}_{k}|:k\in S\}$, as the basis for establishing the results in Theorem 2. For instance, if $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ is bounded away from zero, then it can be verified that $\lambda=\{n/\log(p)\}^{-t}K^{v}$ for some $t\in(1/2,1)$ and $v\in[2t-3/2,2t-1]$ suffices for Condition 4 to hold. If this is relaxed and we permit $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ to tend to zero at a rate satisfying $\min\{|\alpha^{(0)}_{k}|:k\in S\}\{n/\log(p)\}^{m}\to\infty$ for some $m\in(0,1/4]$, then $\lambda=O(\{\log(p)/n\}^{q})$ for some $q\in[m+1/2,3/4]$ will satisfy the requirements. In practice, since $\min\{|\alpha^{(0)}_{k}|:k\in S\}$ is unknown, we adopt a data-driven approach to select $\lambda$ as discussed in the next section.
@@ 4 Simulation Study
We conduct a numerical study to evaluate the finite sample performance of the proposed regularized pseudo-likelihood estimator for the Ising similarity regression model (2.4). Briefly, we consider sample sizes $n\in\{50,100,200,400\}$, numbers of binary responses $p\in\{10,25,50,100,200\}$, and $K=20$ similarity matrices with only the first $K_{0}=5$ of them having non-zero true regression coefficients i.e., $S=\{1,\cdots,5\}$. A total of 1000 datasets are simulated from the Ising similarity regression model (2.4) for each combination of $n$ and $p$. Details of the simulation settings, including the true values of the main effect parameters $\{\theta_{jj}^{(0)}:j=1,\cdots,p\}$ and regression coefficients $\{\alpha^{(0)}_{k}:k=1,\cdots,K\}$ are provided in Section LABEL:sec:simulation_supp of the supplementary material.
@@
We employ a two-step algorithm to compute the regularized pseudo-likelihood estimator. First, from the discussion below equations (2.6) – (2.7), the unregularized estimator is obtained by fitting a logistic regression model with $\bm{y}_{1:n}=(y_{11},\cdots,y_{n1},\cdots,y_{1p},\cdots,y_{np})^{\top}$ as the response and $(\bm{I}_{p}\otimes\bm{1}_{n},\bm{\mathcal{X}})$ as the model matrix, where $\bm{1}_{n}$ is an $n$-dimensional vector of ones, $\otimes$ is the Kronecker product operator, $\bm{\mathcal{X}}=(\bm{\mathcal{X}}^{(1,1)},\cdots,\bm{\mathcal{X}}^{(n,1)},\cdots,\bm{\mathcal{X}}^{(1,p)},\cdots,\bm{\mathcal{X}}^{(n,p)})^{\top}$, and $\bm{\mathcal{X}}^{(i,j)}$ denotes the $j$-th row of the matrix $\bm{\mathcal{X}}^{(i)}$ given in equation (3.9). In the second step, we compute the regularized pseudo-likelihood estimator by fitting an adaptive lasso regularized logistic regression to $\bm{y}_{1:n}$ and $(\bm{I}_{p}\otimes\bm{1}_{n},\bm{\mathcal{X}})$ via the R package glmnet (Friedman, Hastie, and Tibshirani, 2010), where the adaptive regularization weights in (2.7) are constructed based on the unregularized estimator.
To select the tuning parameter $\lambda$ in (2.7), we utilize a ten-fold cross-validation approach, where the observations in each dataset are grouped at the level indexed by $i=1,\cdots,n$. That is, we split the data $\{(\bm{y}_{1},\bm{\mathcal{X}}^{(1)}),\cdots,(\bm{y}_{n},\bm{\mathcal{X}}^{(n)})\}$ randomly into ten folds and select an optimal $\lambda$ that gives the largest mean pseudo-likelihood (averaged across ten test sets), noting that the same set of adaptive weights constructed based on the unregularized estimator from the full dataset is used throughout the cross-validation. The grouping of the observations is done to preserve the dependence structure of the data, since $y_{ij}$ are independent in the index $i$ but dependent in the response index $j$ through the Ising model (see also Warton, Thibaut, and Wang, 2017).
We compare our proposed regularized pseudo-likelihood estimator (Regularized) with three basic approaches: the lasso-regularized estimator (Lasso) based on (2.7) with $w_{k}=1$ for $k=1,\cdots,K$ and $\lambda$ selected using a similar ten-fold cross-validation approach, the unregularized estimator (Unregularized) based on (2.7) with $\lambda=0$, and the oracle estimator (Oracle) based on minimizing an alternative to (2.7) where $\lambda=0$ and $\alpha_{k}$ are set to be zeros for $k\in S^{c}$. The lasso-regularized estimator is considered to study the effect of penalty choice on the empirical performance of the estimator, the unregularized estimator is included to examine whether regularization could lead to better overall estimation performance, and the oracle estimator is used as a benchmark since it incorporates additional information regarding the index set $S=\{1,\cdots,5\}$. We compare point estimation performance for all four estimators of the regression coefficients and main effect parameters using their mean square error (MSE), $\textrm{MSE}_{\alpha}$ and $\textrm{MSE}_{\theta}$, respectively, and the model selection performance of the proposed estimator and lasso-regularized estimator of the regression coefficients using the true positive rate (TPR) and false positive rate (FPR); see Section LABEL:sec:simulation_supp of the supplementary material for details on the computation of these performance measures.